How many planes are determined by three noncollinear?
1 line cause every plane contains atleast 3 or more noncollinear points
A dice is rolled three times the probability of getting at least one odd number is?
Possible outcomes of one roll = 6.
Probability of an even number on one roll = 3/6 = 0.5 .
Probability of an even number on the second roll = 0.5 .
Probability of an even number on the third roll = 0.5 .
Probability of an even number on all three rolls = (0.5 x 0.5 x 0.5) = 0.125 = 1/8
The probability of at least one odd number is the probability of not getting
an even number on all 3 rolls. That's (1 - 1/8) = 7/8 or 0.875 or 87.5% .
Let's consider first one bag:
The probability to grab one green marble is Pg=1/5
The probability to grab one red marble is Pr=1-Pg=4/5
For the 4 bags, it's a binomial distribution: probability to get k green out of n bags,
P(X=k)=nCk pk (1-p)n-k = nCk Pgk Prn-k
Now, the probability to grab at least 2 green marbles from 4 bags is
1 - (Probability to get no green marble from 4 bags) - (Probability to get just one green marble from 4 bags)
Probability to get no green marble is = 4C0 (4/5)4 = (4/5)4
(n=4, k=0, no green marble from each bag, 4C0 Pr4)
Probability to get just one green marble is = 4C1 (1/5) (4/5)3
(n=4, k=1, one green marble from one bag and red marbles from the other ones, 4C1 Pg Pr3)
Probability to grab at least 2 green marbles from 4 bags is
1-(4/5)4-4*(1/5)(4/5)3 = 0.1801
If you were to flip a coin 4 times what is the probability of getting heads at least one time?
As the question is "what is the probability of getting at least one head" the correct way to answer this is to ask what is the probability of not getting any heads and then subtract this from 1.
The probability of not getting a head in 4 flips = 0.54 (i.e. 0.5 * 0.5 * 0.5 * 0.5) = 1/16.
Therefore the probability of getting at least one head is 1 - 1/16 = 15/16.
What is the addition rule of probability?
The addition rule of probability states that the probability that one or the other will happen is the probability of one plus the probability of the other. This rule only applies to mutually exclusive events. For example, the probability that a dice roll will be a 3 is 1/6. The probability that the dice roll will be even is 1/2. These are mutually exclusive events as the dice cannot be both 3 and even. Thus the probability of the dice roll coming up either a 3, or even, is 1/2 + 1/6 = 2/3.
There are 7 balls in total.
4 are red and 1 is the blue "5" ball. Thus there are five balls out of the seven which meet the criteria and therefore the odds must be 5/7.
The probabilities are mutually exclusive (i.e. you can not both draw a red ball and the "5" ball at the same time as the "5" is blue) so you just add the probabilities of each together.
Probability of drawing a red ball = 4/7 (4 of the 7 balls are red)
Probability of drawing the "5" ball = 1/7 (1 of the 7 balls is the "5" ball)
Overall probability = 4/7 + 1/7 = 5/7
How do you calculate standard deviation with the help of z-score?
A z-score cannot help calculate standard deviation. In fact the very point of z-scores is to remove any contribution from the mean or standard deviation.
What is the probability of not rolling a sum of 7 three times in a row?
Possible outcomes for one cube = 6 .
Possible outcomes for 2 cubes = 6 x 6 = 36 .
Ways to roll a sum of 7 = 6 ways,
1, 6 . . . 2, 5 . . . 3, 4
6, 1 . . . 5, 2 . . . 4, 3
Probability of rolling a 7 on the first roll = 6/36 = 1/6 .
Probability of rolling a 7 on the second roll = 6/36 = 1/6 .
Probability of rolling a 7 on the third roll = 6/36 = 1/6 .
Probability of rolling a sum of 7 three times in a row = (1/6 x 1/6 x 1/6) = (1/6)3 = 1/216 = 0.00463 = 0.463%
Probability of not rolling a sum of 7 three times in a row = 1 - (probability of doing it) = 99.537%
How do you do the chapter 4 frappy?
read chapter 4 in your textbook, take notes, practice some problems an then try to do the frappy
What is the probability of picking an ace from a pack of playing cards?
The probability of drawing an Ace from a standard deck of 52 cards is 4 in 52, or 1 in 13, or about 0.07692.
What is the probability of getting court card?
The probability of drawing a court card (a jack, queen, or king) from a standard deck of 52 cards is 12 in 52, or 3 in 13, or about 0.2308.
If you count the ace as a court card, then the probability is 16 in 52, or 4 in 13, or about 0.3077.
A coin is tossed 4 times. What is the probability of getting tails 4 times in a row?
The probability to get tails once is 1/2 (for a fair coin)
The probability to get tails twice = the probability to get it once x the probability to get it a second time
The probability to get tails 4 times in a row is (1/2)4=1/16
The probability to get tails n times in a row is (1/2)n=1/2n
The same thing is also true for heads (same probability: 1/2 each time)
Contoh jawapan kerja projek matematik tambahan perak 2011?
Bincangkan secara ringkas bagaimana Statistik boleh digunakan dalam kehidupan harian.Sertakan gambar yang sesuai
What is the probability of getting both tails on two tosses of a coin?
The probability of two tails on two tosses of a coin is 0.52, or 0.25.
How do you calculate the standard deviation of the mean using Excel?
=stdev(...) will return the N-1 weighted sample standard deviation.
=stdevp(...) will return the N weighted population standard deviation.
What is the probability of rolling a sum of 8 and doubles when rolling two dice?
The probability of rolling a sum of 8 and doubles when rolling two dice is 1 in 36, or about 0.02778.
Simply note that there are 36 permutations of two dice, of which exactly one of them (a 4 and a 4) matches the conditions specified.
How many 6 of diamonds in a deck of 52 cards?
There is only 1 in each deck.
There are 4 '6s' in total:
- 6 of hearts
- 6 of clubs
- 6 of spades
- 6 of diamonds
What does it mean when a data set of 10 measurements has a mean of 15 and a standard deviation of 0?
It means that all of the ten numbers are 15!
Standard deviation tells you how spread out the data is from the mean value. Or in other words, it tells you how far the numbers in your data are away from the mean value.
If the standard deviation is a high number, it means the data is largely spread out and that there are big differences in the data. The numbers in the data would be quite far from each other. For example, if you had data like: 8, 35, 13, 47, 22, 64, this would probably mean that you'll get a high standard deviation because each of the numbers are very spread out.
On the other hand, if the standard deviation is small, it tells you that the numbers in the data are quite close together and that there is only a small difference between the numbers in the data. For example, if you had data like: 19, 25, 20, 22, 23, 18, this would probably mean that you'll get a low standard deviation because each of the numbers aren't that spread out
In the scenario you've given, the standard deviation is ZERO. This means that there is no spread or variation AT ALL with the numbers in your data. This means every single number in the data is the same.
Since your mean is 15 and every number in your data is the same, that means that all the ten numbers in your data have to be 15!
Hope that makes sense.
Jamz159
What is the probability of a number cube?
A number cube has six faces, so the probability of any one of them showing on a fair throw is 1 in 6, or about 0.1667.
68 % is about one standard deviation - so there score would be between 64 and 80 (72 +/- 8)
What is difference between sample and representative sample?
A sample is any subset of the total population. A representative sample is one that is chosen so that its characteristics are similar to that of the population.
Why is heads more likely than tails?
It isnt, There is a 50/50 chance on getting a head or either a tail, as there are 2 sides as it has 2 faces. I hope this helped :D X